# RoboRefer: Towards Spatial Referring with Reasoning in Vision-Language Models for Robotics

### v3

@@ 1 Introduction
Spatial referring task comprises two distinct levels of complexity: (1) Single-step spatial understanding, which forms the foundation of spatial perception by accurately recognizing objects’ spatial properties (e.g., position, orientation) and their spatial relations (e.g., distance, direction). This level, where most current research [1, 2, 3, 4, 5, 6, 7] concentrates, provides the essential perceptual basis for complex spatial referring. (2) Multi-step spatial reasoning, which transcends basic understanding through compositional reasoning to resolve complex spatial references sequentially. Despite its importance for sophisticated spatial intelligence, this capability remains underexplored. Thus, this work attempts to address this gap by integrating both levels for comprehensive spatial referring. In Fig 1, one must first identify the plate closest to the observer and locate the desired soy sauce dish, then determine the free space between them, which is increasingly challenging as more spatial constraints are introduced.
Specifically, existing vision-language models (VLMs) [8, 9, 10, 11]-based methods mainly attempt to enhance the first level, i.e., single-step spatial understanding by integrating 3D inputs. However, they either demand costly 3D reconstruction of multi-view images [12, 13], causing modality gaps , or treat depth as RGB-like inputs [1, 14, 3] via a shared image encoder, risking modality interference and degrading pretrained image encoders, requiring additional co-training data for compensation. In contrast, the second level, i.e., multi-step spatial referring with reasoning, remains underexplored due to the scarcity of suitable datasets, limiting current models’ capability and preventing exploration of how single-step understanding might support it. Moreover, current VLMs depend heavily on supervised fine-tuning (SFT) for implicit reasoning, risking memorizing answers over explicit reasoning and thereby hindering generalization and accuracy in open-world spatial referring.
@@
To advance spatial referring, we introduce RefSpatial, a large-scale dataset of 2.52.5M high-quality examples with 2020M QA pairs ($2\times$ prior [3]). Leveraging diverse data sources from 2D/3D/Simulation, this dataset can teach a general VLM to achieve spatial referring in a bottom-up manner. Specifically, 2D web images provide fundamental spatial concepts and broad depth perception (indoor and outdoor), 3D embodied videos refine fine-grained spatial understanding of indoor scenes for robotics, and simulated data with ground-truth reasoning processes encourage multi-step spatial referring (up to $5$ steps). Notably, RefSpatial includes $31$ spatial relations, far exceeding $15$ found in previous datasets [2, 3], and each sample contains RGB-D data to support depth alignment in SFT stage.
We evaluate our SFT-trained model on existing single-step spatial reasoning benchmarks (e.g., CV-Bench [15], BLINK [16]), achieving SOTA performance with an average success rate of 89.689.6%. To address the lack of multi-step spatial referring benchmarks, we introduce RefSpatial-Bench, comprising $200$ real-world images with manually annotated tasks for object location and placement. Over 7070% of the samples require multi-step reasoning (up to $5$ steps) and are annotated with precise masks. Our model consistently outperforms all baselines on this benchmark, even surpassing Gemini-2.5-Pro by an average of 17.417.4%. Moreover, in Fig. 1 and Sec. 4.4, RoboRefer can execute long-horizon, dynamic tasks in cluttered real-world scenes with various control policies, exhibiting strong generalization across robots (e.g., UR5, G1 humanoid) and tasks (e.g., manipulation, navigation).
@@ 2 Related work
Spatial Understanding with VLMs. Spatial understanding [17, 16, 18, 19, 20, 21, 22, 23] focuses on object-centric properties (e.g., position, orientation) and inter-object relations (e.g., distance, direction), while spatial reasoning [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35] draws higher-level inferences over such information. Recent advances in VLMs [8, 9, 36, 37, 11, 10, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56] enhance these two abilities via two paradigms: (1) tool-based approaches [57, 7, 14, 58, 59, 60, 61, 62, 63] that integrate VLMs with vision foundation models [64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77] to extract and reason spatial cues and (2) data-driven methods, which fine-tune VLMs using pseudo-3D annotations [6, 1], real-world 3D datasets [2, 3], or simulated data [4, 78]. However, existing datasets lack multi-step reasoning annotations critical for spatial referring tasks, and a benchmark for evaluating such abilities remains unavailable. We thus introduce a new dataset and benchmark specifically tailored for spatial referring.
@@
Referring with VLMs for Robotics. Referring, also known as Referring Expression Comprehension (REC) [79, 80, 81, 82, 83, 84, 85, 86], leverages unambiguous descriptions to localize a unique region/point in an image, and has seen great progress via VLMs [87, 88, 89, 90, 91, 92]. Unlike Phrase Localization [93, 94, 95] and Generalized Visual Grounding [96, 97, 98, 99, 100], which address ambiguous or multiple referents, REC focuses on one single target—an emphasis crucial for robotics, especially in pick-and-place tasks requiring precise object identification and destination [101, 102, 103, 104, 105]. While 2D REC relies on object attributes (e.g., color) and image-plane localization (e.g., top right of the image), real-world scenarios for robotics require 3D spatial reasoning to localize (e.g., “near” vs. “far”). Although efforts [106, 107, 108] like RoboPoint [5] incorporate basic spatial cues via images to meet such expectations, they often struggle with complex environments and instructions required for spatial referring. Thus, we propose RoboRefer, a 3D-aware framework that employs multi-step reasoning to ensure precise spatial referring for robotics.
@@
Reinforcement Fine-tuning for VLMs. Reinforcement Fine-tuning (RFT) [109, 110, 111, 112, 113] is a post-training strategy that aligns models with human preferences or specific goals via feedback, complementing SFT [114, 115], which adapts pre-trained models using task-oriented data. Recent advances in LLM-based reasoning [116, 117, 113, 118, 119] have shifted RL in VLMs toward visual reasoning [120, 121, 122, 123, 124], grounding [125, 126, 127], segmentation [128] and trajectory prediction [129]. However, most methods rely solely on 2D perception, limiting their ability to handle spatial referring tasks that require 3D spatial reasoning. To address this, we propose a two-stage training strategy: (1) incorporate depth information during SFT to strengthen spatial understanding; (2) RFT stage then leverages intermediate perception outputs powered by SFT to enable multi-step spatial referring with reasoning.
@@ 3.1 Problem Formulation
Unlike region-based 2D referring methods [100, 87, 92], this point-based formulation is more suitable and generalizable for robotics. Compared to the 2D bbox, point can naturally map to 3D coordinates via depth, providing accurate spatial anchors. By leveraging predicted points for navigation, grasping, or placement, this formulation enables multi-task learning and execution. Moreover, it can accurately localize a visible object part in occlusion scenarios, while 2D bbox often includes irrelevant objects.
@@ 3.2 RoboRefer: A 3D-aware reasoning VLM for spatial referring
VLM Architecture. In Fig. 2, RoboRefer employs separate RGB and depth encoders to extract features, which are then aligned via projectors with the LLM for VQA or point prediction. As 3D cues are vital for spatial understanding, 2D VLMs pretrained solely on RGB lack accurate 3D perception. Recent methods [1, 3, 14] avoid explicit 3D representations by treating depth as an image-like modality and sharing the RGB encoder, but this causes modality interference, degrading the pretrained encoder and requiring additional RGB co-training to compensate. To address this, we propose a simple yet effective approach: a dedicated depth encoder and projector, initialized from their RGB counterparts. Notably, during joint RGB and RGB-D training, the image encoder remains unaffected by depth input, while the depth encoder is updated independently. This design not only avoids modality interference and preserves general VQA performance without extensive RGB-only co-training, but it also improves spatial understanding through enhanced perception of depth cues (e.g., distance, near–far relations, and perspective-based size variations). See Appx. D.1 for details.
@@
Supervised Fine-tuning. We adopt NVILA [37] as our base VLM; however, its 2D-only pretraining limits spatial understanding. To address this, we propose a two-step SFT. (1) Depth alignment. In Fig. 2, we first train a depth projector to align the newly introduced depth space with the textual space, leveraging RGB-D annotations of the RefSpatial (see Sec. 3.3). In this step, only the depth projector is updated. (2) Spatial understanding enhancement. We fine-tune all parameters on the RefSpatial, including single-step fine-grained annotations and multi-step reasoning data with explicit reasoning processes, and additional instruction-following datasets [130, 131, 86]. Therefore, the model is jointly optimized on RGB and RGB-D inputs, with separate updates for the image and depth encoders. This process not only enhances single-step spatial understanding via the new depth modality but also bolsters implicit multi-step reasoning through data with explicit reasoning processes, providing a “cold start” for the subsequent RFT stage. As a result, this SFT-trained model exhibits improved capability for multi-step spatial referring tasks. Please check Appx. D.3 for details.
@@
Reinforcement Fine-tuning. Though SFT employs data with precise reasoning, it tends to memorize answers rather than generalize to novel spatial constraints. We thus design a subsequent RFT stage using Group Relative Policy Optimization (GRPO [113]) with multi-step reasoning data from RefSpatial. To guide RFT for more accurate point predictions, we first define two outcome reward (i.e., only care about whether the output answer is correct) functions: (1) Outcome Format Reward ($R_{OF}$) for structured reasoning and clarity; and (2) Point L1 Reward ($R_{P}$) granting a score of $1$ if the final prediction falls within a specific range near the ground-truth point, and 0 otherwise. To enhance intermediate reasoning precision, we exploit key-step perception annotations from RefSpatial and design specialized metric-sensitive process reward functions: (1) Process Format Reward ($R_{PF}$), enforcing the format “[Perception Type] [Target Object]:”; (2) Accuracy Reward ($R_{Acc}$), which applies to steps included in the key-step perception annotations. For each relevant step, we measure the prediction error using a specific metric, according to the perception type (e.g., L1 distance for positions between ground-truth points and predicted points). Notably, this design is order-invariant and does not constrain the reasoning trajectory to a fixed sequence. We sample $N$ responses $\{a_{1},\ldots,a_{N}\}$ from the current policy (initialized from the SFT model) to encourage exploration. Each response receives a combined reward ($r_{i}=R_{OF}(a_{i})+R_{P}(a_{i})+\alpha R_{PF}(a_{i})+\alpha R_{Acc}(a_{i})$). where $\alpha$ is set to 0.25. Rewards are normalized within each group to compute relative advantages ($A_{i}=\frac{r_{i}-\text{mean}(\{r_{j}\})}{\text{std}(\{r_{j}\})}$), which are then used to update the policy, reinforcing high-quality responses and suppressing suboptimal ones. A KL-divergence regularization term stabilizes updates by constraining them near the reference policy. Notably, the SFT initialization provides a strong prior, enabling rapid adaptation to output formats and supporting accurate, step-wise spatial reasoning by using the spatial understanding learned from SFT. Fig 2 shows that the RFT-trained model generalizes well to tasks like $4$-step spatial referring, progressively handling intricate spatial relations, and yielding precise point predictions. For more details about the RFT training and reward design, please see Appx. D.4.
@@ 3.3.1 Overview
RefSpatial is a comprehensive dataset integrating 2D images from OpenImages [132], 3D embodied videos from CA-1M [133], and simulated scenes from Infinigen [134] using Objaverse [135] assets (See Fig. 3 (a)). RefSpatial’s key features are: (1) Fine-Grained Annotations. While prior spatial datasets [3, 2] simplify object reference by limiting each category to a single instance per scene, RefSpatial includes multiple objects of the same category. Moreover, each object is annotated with hierarchical captions—from broad categories (e.g., “cup”) to precise spatial referents (e.g., “the third cup from the left”, “the cup closest to the camera”)—enabling unambiguous spatial referring in cluttered environments. (2) Multi-Dimensionality. Beyond basic spatial concepts, relations, point coordinates, and point depth predictions, the dataset supports multi-step spatial reasoning by annotating detailed reasoning processes (all simulated data), addressing limitations in existing datasets. (3) High Quality. We rigorously filter data to maintain quality. Retain 466k OpenImages containing text-referable, spatially relevant objects (down from 1.7M); sample 100k frames from CA-1M with text-identifiable 3D bounding boxes (down from 2M); and manually check and annotate 3k Objaverse-LVIS assets with semantic orientation labels (down from 46k). (4) Large Scale. Comprising 2.5M samples and 20M QA pairs, our dataset spans qualitative VQA, quantitative queries on object attributes/relations, and point coordinate prediction (Fig. 3(b)). (5) Rich Diversity. RefSpatial spans indoor and outdoor scenes, covers common embodied scenes and integrates $31$ distinct spatial relations (See Fig. 3 (c)), fostering precise spatial understanding during SFT. (6) Easy Scalability. Our pipeline seamlessly scales spatial referring data using diverse sources, including 2D images, 3D videos with bounding boxes, and simulation assets. See Appx. B for more dataset details.

### v4

@@ 1 Introduction
Spatial referring task comprises two distinct levels of complexity: (1) Single-step spatial understanding, which forms the foundation of spatial perception by accurately recognizing objects’ spatial properties (e.g., position, orientation) and their spatial relations (e.g., distance, direction). This level, where most current research [cheng2024spatialrgpt, song2024robospatial, daxberger2025mm, ray2024sat, yuan2024robopoint, chen2024spatialvlm, qi2025sofar] concentrates, provides the essential perceptual basis for complex spatial referring. (2) Multi-step spatial reasoning, which transcends basic understanding through compositional reasoning to resolve complex spatial references sequentially. Despite its importance for sophisticated spatial intelligence, this capability remains underexplored. Thus, this work attempts to address this gap by integrating both levels for comprehensive spatial referring. In Fig 1, one must first identify the plate closest to the observer and locate the desired soy sauce dish, then determine the free space between them, which is increasingly challenging as more spatial constraints are introduced.
Specifically, existing vision-language models (VLMs) [achiam2023gpt, team2023gemini, deitke2024molmo, bai2025qwen2]-based methods mainly attempt to enhance the first level, i.e., single-step spatial understanding by integrating 3D inputs. However, they either demand costly 3D reconstruction of multi-view images [hong20233dllm, hong20233d], causing modality gaps , or treat depth as RGB-like inputs [cheng2024spatialrgpt, cai2024spatialbot, daxberger2025mm] via a shared image encoder, risking modality interference and degrading pretrained image encoders, requiring additional co-training data for compensation. In contrast, the second level, i.e., multi-step spatial referring with reasoning, remains underexplored due to the scarcity of suitable datasets, limiting current models’ capability and preventing exploration of how single-step understanding might support it. Moreover, current VLMs depend heavily on supervised fine-tuning (SFT) for implicit reasoning, risking memorizing answers over explicit reasoning and thereby hindering generalization and accuracy in open-world spatial referring.
@@
To advance spatial referring, we introduce RefSpatial, a large-scale dataset of 2.52.5M high-quality examples with 2020M QA pairs ($2\times$ prior [daxberger2025mm]). Leveraging diverse data sources from 2D/3D/Simulation, this dataset can teach a general VLM to achieve spatial referring in a bottom-up manner. Specifically, 2D web images provide fundamental spatial concepts and broad depth perception (indoor and outdoor), 3D embodied videos refine fine-grained spatial understanding of indoor scenes for robotics, and simulated data with ground-truth reasoning processes encourage multi-step spatial referring (up to $5$ steps). Notably, RefSpatial includes $31$ spatial relations, far exceeding $15$ found in previous datasets [song2024robospatial, daxberger2025mm], and each sample contains RGB-D data to support depth alignment in SFT stage.
We evaluate our SFT-trained model on existing single-step spatial reasoning benchmarks (e.g., CV-Bench [tong2024cambrian], BLINK [fu2024blink]), achieving SOTA performance with an average success rate of 89.689.6%. To address the lack of multi-step spatial referring benchmarks, we introduce RefSpatial-Bench, comprising $200$ real-world images with manually annotated tasks for object location and placement. Over 7070% of the samples require multi-step reasoning (up to $5$ steps) and are annotated with precise masks. Our model consistently outperforms all baselines on this benchmark, even surpassing Gemini-2.5-Pro by an average of 17.417.4%. Moreover, in Fig. 1 and Sec. 4.4, RoboRefer can execute long-horizon, dynamic tasks in cluttered real-world scenes with various control policies, exhibiting strong generalization across robots (e.g., UR5, G1 humanoid) and tasks (e.g., manipulation, navigation).
@@ 2 Related work
Spatial Understanding with VLMs. Spatial understanding [azuma2022scanqa, fu2024blink, hudson2019gqa, jia2024sceneverse, johnson2017clevr, szymanska2024space3d, du2024embspatial, ramakrishnan2024does, zhou2025robotracer] focuses on object-centric properties (e.g., position, orientation) and inter-object relations (e.g., distance, direction), while spatial reasoning [kamath2023s, liu2023visual, rajabi2023towards, ranasinghe2024learning, shiri2024empirical, lee2025perspective, wang2024picture, tang2024sparkle, liu2025spatialcot, liu2025ssr, ma2025spatialllm, ma2025spatialreasoner] draws higher-level inferences over such information. Recent advances in VLMs [achiam2023gpt, team2023gemini, anthropic2024claude, liu2024nvila, bai2025qwen2, deitke2024molmo, yin2023lamm, qin2024mp5, qin2025navigatediff, qin2024worldsimbench, chen2024rh20t, zhou2024minedreamer, ji2025robobrain, huang2024story3d, li2025t2isafety, zhu2025internvl3, zhang2024navid, zhang2024uninavid, liu2024robomamba, team2025robobrain, ji2025mathsticks, zhang2025beyond, li2025labutopia, wang2025trackvla, liu2023visual_instruction_tuning] enhance these two abilities via two paradigms: (1) tool-based approaches [maspatialpin, qi2025sofar, cai2024spatialbot, ding2024open6dor, qin2025robofactory, zhu2025dexflywheel, tan2025roboos, zhou2024code, han2025tiger] that integrate VLMs with vision foundation models [jin2023perspective, zhang2024recognize, liu2024segment, piccinelli2025unidepthv2, bochkovskii2024depth, yuan2025motiontrans, li2024lamp, li2024mulsmo, li2025languagelocomotionretargetingfreehumanoid, tschannen2025siglip, zhai2023sigmoid, zhu2023tame, kirillov2023segment, ravi2024sam] to extract and reason spatial cues and (2) data-driven methods, which fine-tune VLMs using pseudo-3D annotations [chen2024spatialvlm, cheng2024spatialrgpt], real-world 3D datasets [song2024robospatial, daxberger2025mm], or simulated data [ray2024sat, wang2025pulsecheck457]. However, existing datasets lack multi-step reasoning annotations critical for spatial referring tasks, and a benchmark for evaluating such abilities remains unavailable. We thus introduce a new dataset and benchmark specifically tailored for spatial referring.
@@
Referring with VLMs for Robotics. Referring, also known as Referring Expression Comprehension (REC) [mao2016generation, van2006building, golland2010game, mitchell2010natural, mitchell2013generating, fitzgerald2013learning, kazemzadeh2014referitgame, yu2016modeling], leverages unambiguous descriptions to localize a unique region/point in an image, and has seen great progress via VLMs [you2023ferret, peng2023kosmos, chen2023shikra, zhao2023bubogpt, wang2023visionllm, xiao2024florence]. Unlike Phrase Localization [plummer2015flickr30k, wang2016structured, plummer2017phrase] and Generalized Visual Grounding [xie2023described, he2023grec, liu2023gres, hu2023beyond, liu2024grounding], which address ambiguous or multiple referents, REC focuses on one single target—an emphasis crucial for robotics, especially in pick-and-place tasks requiring precise object identification and destination [zeng2021transporter, liu2022structformer, liu2023structdiffusion, yuan2023m2t2, shridhar2022cliport]. While 2D REC relies on object attributes (e.g., color) and image-plane localization (e.g., top right of the image), real-world scenarios for robotics require 3D spatial reasoning to localize (e.g., “near” vs. “far”). Although efforts [team2025gemini, nasiriany2024pivot, yang2025magma] like RoboPoint [yuan2024robopoint] incorporate basic spatial cues via images to meet such expectations, they often struggle with complex environments and instructions required for spatial referring. Thus, we propose RoboRefer, a 3D-aware framework that employs multi-step reasoning to ensure precise spatial referring for robotics.
@@
Reinforcement Fine-tuning for VLMs. Reinforcement Fine-tuning (RFT) [bai2022training, bai2022constitutional, rafailov2023direct, an2025agfsync, shao2024deepseekmath] is a post-training strategy that aligns models with human preferences or specific goals via feedback, complementing SFT [wei2021finetuned, zhou2023lima], which adapts pre-trained models using task-oriented data. Recent advances in LLM-based reasoning [jaech2024openai, guo2025deepseek, shao2024deepseekmath, zhang2025landscape, zhou2025reso] have shifted RL in VLMs toward visual reasoning [yang2025r1, zhang2025r1, zhang2025mm, tan2025reason, kang2025viki, tan2025robo, zhang2025prune4web], grounding [liu2025visual, zhan2025vision, shen2025vlm], segmentation [liu2025seg] and trajectory prediction [song2025maniplvm]. However, most methods rely solely on 2D perception, limiting their ability to handle spatial referring tasks that require 3D spatial reasoning. To address this, we propose a two-stage training strategy: (1) incorporate depth information during SFT to strengthen spatial understanding; (2) RFT stage then leverages intermediate perception outputs powered by SFT to enable multi-step spatial referring with reasoning.
@@ 3.1 Problem Formulation
Unlike region-based 2D referring methods [liu2024grounding, you2023ferret, xiao2024florence], this point-based formulation is more suitable and generalizable for robotics. Compared to the 2D bbox, point can naturally map to 3D coordinates via depth, providing accurate spatial anchors. By leveraging predicted points for navigation, grasping, or placement, this formulation enables multi-task learning and execution. Moreover, it can accurately localize a visible object part in occlusion scenarios, while 2D bbox often includes irrelevant objects.
@@ 3.2 RoboRefer: A 3D-aware reasoning VLM for spatial referring
VLM Architecture. In Fig. 2, RoboRefer employs separate RGB and depth encoders to extract features, which are then aligned via projectors with the LLM for VQA or point prediction. As 3D cues are vital for spatial understanding, 2D VLMs pretrained solely on RGB lack accurate 3D perception. Recent methods [cheng2024spatialrgpt, daxberger2025mm, cai2024spatialbot] avoid explicit 3D representations by treating depth as an image-like modality and sharing the RGB encoder, but this causes modality interference, degrading the pretrained encoder and requiring additional RGB co-training to compensate. To address this, we propose a simple yet effective approach: a dedicated depth encoder and projector, initialized from their RGB counterparts. Notably, during joint RGB and RGB-D training, the image encoder remains unaffected by depth input, while the depth encoder is updated independently. This design not only avoids modality interference and preserves general VQA performance without extensive RGB-only co-training, but it also improves spatial understanding through enhanced perception of depth cues (e.g., distance, near–far relations, and perspective-based size variations). See Appx. D.1 for details.
@@
Supervised Fine-tuning. We adopt NVILA [liu2024nvila] as our base VLM; however, its 2D-only pretraining limits spatial understanding. To address this, we propose a two-step SFT. (1) Depth alignment. In Fig. 2, we first train a depth projector to align the newly introduced depth space with the textual space, leveraging RGB-D annotations of the RefSpatial (see Sec. 3.3). In this step, only the depth projector is updated. (2) Spatial understanding enhancement. We fine-tune all parameters on the RefSpatial, including single-step fine-grained annotations and multi-step reasoning data with explicit reasoning processes, and additional instruction-following datasets [liu2024mitigating, liu2024improved, yu2016modeling]. Therefore, the model is jointly optimized on RGB and RGB-D inputs, with separate updates for the image and depth encoders. This process not only enhances single-step spatial understanding via the new depth modality but also bolsters implicit multi-step reasoning through data with explicit reasoning processes, providing a “cold start” for the subsequent RFT stage. As a result, this SFT-trained model exhibits improved capability for multi-step spatial referring tasks. Please check Appx. D.3 for details.
@@
Reinforcement Fine-tuning. Though SFT employs data with precise reasoning, it tends to memorize answers rather than generalize to novel spatial constraints. We thus design a subsequent RFT stage using Group Relative Policy Optimization (GRPO [shao2024deepseekmath]) with multi-step reasoning data from RefSpatial. To guide RFT for more accurate point predictions, we first define two outcome reward (i.e., only care about whether the output answer is correct) functions: (1) Outcome Format Reward ($R_{OF}$) for structured reasoning and clarity; and (2) Point L1 Reward ($R_{P}$) granting a score of $1$ if the final prediction falls within a specific range near the ground-truth point, and 0 otherwise. To enhance intermediate reasoning precision, we exploit key-step perception annotations from RefSpatial and design specialized metric-sensitive process reward functions: (1) Process Format Reward ($R_{PF}$), enforcing the format “[Perception Type] [Target Object]:”; (2) Accuracy Reward ($R_{Acc}$), which applies to steps included in the key-step perception annotations. For each relevant step, we measure the prediction error using a specific metric, according to the perception type (e.g., L1 distance for positions between ground-truth points and predicted points). Notably, this design is order-invariant and does not constrain the reasoning trajectory to a fixed sequence. We sample $N$ responses $\{a_{1},\ldots,a_{N}\}$ from the current policy (initialized from the SFT model) to encourage exploration. Each response receives a combined reward ($r_{i}=R_{OF}(a_{i})+R_{P}(a_{i})+\alpha R_{PF}(a_{i})+\alpha R_{Acc}(a_{i})$). where $\alpha$ is set to 0.25. Rewards are normalized within each group to compute relative advantages ($A_{i}=\frac{r_{i}-\text{mean}(\{r_{j}\})}{\text{std}(\{r_{j}\})}$), which are then used to update the policy, reinforcing high-quality responses and suppressing suboptimal ones. A KL-divergence regularization term stabilizes updates by constraining them near the reference policy. Notably, the SFT initialization provides a strong prior, enabling rapid adaptation to output formats and supporting accurate, step-wise spatial reasoning by using the spatial understanding learned from SFT. Fig 2 shows that the RFT-trained model generalizes well to tasks like $4$-step spatial referring, progressively handling intricate spatial relations, and yielding precise point predictions. For more details about the RFT training and reward design, please see Appx. D.4.
@@ 3.3.1 Overview
RefSpatial is a comprehensive dataset integrating 2D images from OpenImages [kuznetsova2020open], 3D embodied videos from CA-1M [lazarow2024cubify], and simulated scenes from Infinigen [raistrick2024infinigen] using Objaverse [deitke2023objaverse] assets (See Fig. 3 (a)). RefSpatial’s key features are: (1) Fine-Grained Annotations. While prior spatial datasets [daxberger2025mm, song2024robospatial] simplify object reference by limiting each category to a single instance per scene, RefSpatial includes multiple objects of the same category. Moreover, each object is annotated with hierarchical captions—from broad categories (e.g., “cup”) to precise spatial referents (e.g., “the third cup from the left”, “the cup closest to the camera”)—enabling unambiguous spatial referring in cluttered environments. (2) Multi-Dimensionality. Beyond basic spatial concepts, relations, point coordinates, and point depth predictions, the dataset supports multi-step spatial reasoning by annotating detailed reasoning processes (all simulated data), addressing limitations in existing datasets. (3) High Quality. We rigorously filter data to maintain quality. Retain 466k OpenImages containing text-referable, spatially relevant objects (down from 1.7M); sample 100k frames from CA-1M with text-identifiable 3D bounding boxes (down from 2M); and manually check and annotate 3k Objaverse-LVIS assets with semantic orientation labels (down from 46k). (4) Large Scale. Comprising 2.5M samples and 20M QA pairs, our dataset spans qualitative VQA, quantitative queries on object attributes/relations, and point coordinate prediction (Fig. 3(b)). (5) Rich Diversity. RefSpatial spans indoor and outdoor scenes, covers common embodied scenes and integrates $31$ distinct spatial relations (See Fig. 3 (c)), fostering precise spatial understanding during SFT. (6) Easy Scalability. Our pipeline seamlessly scales spatial referring data using diverse sources, including 2D images, 3D videos with bounding boxes, and simulation assets. See Appx. B for more dataset details.